OpenStudy (anonymous):

Find all solutions of the following linear system: 2x+4y=2 3x+6y=3

OpenStudy (anonymous):

When I'm trying to solve it I get the two equations are the same, 1x +2y = 1 1x +2y = 1 I'm not sure where to go from here, or wha this means

OpenStudy (anonymous):

y = (1-x)/2?

myininaya (myininaya):

that means the lines are lying on top of each other and you can say they intersect for every point so there are infinitely many solutions

OpenStudy (anonymous):

Thanks so much! that makes sense

myininaya (myininaya):

and you can say they intersect for every point (x,y)* and you are welcome

OpenStudy (anonymous):

Thanks again, I can put together that they intersect for every point [x, (1-x)/2 ]

myininaya (myininaya):

yes that will be beautiful

myininaya (myininaya):

( ) instead of [ ]